# grenedalf: population genetic statistics for the next generation of pool sequencing

### v2

@@ 3 Features and Data Processing Flow
Finally, with the data stream processed as described above, the desired statistical estimators are computed. Typically, the statistics are then averaged over the window, in order to obtain per-base-pair estimates. To account for the characteristics of the input data (with missing data; only containing variant loci; etc), we offer different policies for the window averaging: (a) the window size (likely an underestimation), (b) the number of all available loci in the input, (c) the number of “valid” high-quality positions (i.e., the number of loci that passed all filters; invariants and SNPs), (d) the number of SNPs only (likely an overestimation), and (e) no averaging (i.e., simply report the sum of all per-site values; this allows the user to apply custom averaging later on). These policies are covering the most common use cases of data types. When the data has sufficient coverage, we recommend to use (c), the number of high-quality positions. If specified, this also takes the mask into account, so that SNP-only input data can be properly normalized per window.

### v3

@@ 3 Features and Data Processing Flow
Finally, with the data stream processed as described above, the desired statistical estimators are computed. Typically, the statistics are then averaged over the window, in order to obtain per-base-pair estimates. To account for the characteristics of the input data (with missing data; only containing variant loci; etc), we offer different policies for the window averaging: (a) the window size (likely an underestimation), (b) the number of all available loci in the input, (c) the number of “valid” high-quality positions (i.e., the number of loci that passed all quality filters; invariants and SNPs), (d) the number of SNPs only (likely an overestimation), and (e) no averaging (i.e., simply report the sum of all per-site values; this allows the user to apply custom averaging later on). These policies are covering the most common use cases of data types. When the data has sufficient coverage, we recommend to use (c), the number of high-quality positions. If specified, this also takes the mask into account, so that SNP-only input data can be properly normalized per window.